# Multifractaility, topology and anomalous Hall conductivity on a 30 degrees twisted bilayer honeycomb lattice

### v2

@@ II.2.1 Topological entanglement entropy
Figure 7: A scheme of partitioning the system, which we use to compute the topological contribution $\gamma$ to the entanglement entropy. We fix the external diameter of the susbystems to be $L/2$ and internal diameter to be $L/4$.
We plot the results for $\gamma$ as functions of $m$, $t_{inter}$ on the Figs. 8(a) (we assume $E_{F}=0$). We can see that in the ’weakly coupled topological’ phase, $\gamma$ is a non-zero constant, which in turn confirms that it is a topological quasicrystal. In a similar way, in the ’weakly coupled non-topological’ phase, $\gamma$ is equal to zero, which means that the system is non-topological. In the strongly coupled phase, the concept of topological entanglement entropy becomes ill-defined because the system is gapless.
@@
Figure 8: Topological entanglement entropy $\gamma$ plotted against $t_{inter}$ (8(a)) for $m=0,2$ and against $m$ (8(b)) for $t_{inter}=0.0,1.0,10.0$ It is equal to a non-zero constant in the topological gapped phase and zero in the non-topological phase
II.2.2 Local Chern marker
@@
Figure 9: 9(a) Local Chern marker plotted for lattices with $-L<x,y<L$ at various values of $m$, $t_{inter}$. One can see that in the bulk of a topological phase, it is approximately equal to Chern invariant, and in the non-topological phase, it is equal to zero. In the multifractal phase, it fluctuates within the lattice. 9(b) Averaged local Chern marker $C_{ave}$ and its standard deviation $\Delta C$ taken over the inner part of the lattice with $-L/2<x,y<L/2$. It remains constant and approximately equal to the bulk Chern invariant in the weakly coupled phase, but starts fluctuating at large $t_{inter}$.
II.2.3 Anomalous Hall conductivity
@@
×$\times\left\{\psi^{\dagger}_{e,i}H_{ij}(x_{j}-x_{i})\psi_{f,j}\cdot\psi^{\dagger}_{f,k}H_{kl}(y_{l}-y_{k})\psi_{e,l}\right.$
@@
Let us emphasize once again, that in the Eq. (II.2.3), $\sigma_{ij}$ describes electric current through a bond between the lattice sites $i,j$ (we assumed that the electric field is uniform). However, one can sum electric currents through all bonds next to a given lattice site $i$ and consider $\sigma_{i}=\sum\limits_{j}\sigma_{ij}$. For a crystalline system, one can take a step further and split the lattice site index $i$ into two indices $i^{\prime},\alpha$ numerating unit cells and sublattice degrees of freedom respectively. The well-known integer value of the anomalous Hall conductivity is obtained after an additional summation over $\alpha$. A simple explanation of the latter is as follows: for an infinite crystalline system one may transform any dependency over unit cell index $i^{\prime}$ into momentum representation, and the well-known integer AHE occurs at zero momentum. However, its expression in terms of Berry curvature still contains summation over sublattice degrees of freedom. Nevertheless, on a honeycomb lattice, the two sublattice degrees of freedom are equivalent to each other, and therefore the AHE for each of them would be the same, i.e. half-integer.
More interestingly, properties of AHE obtained from the Eq. (II.2.3) on a finite lattice are fundamentally different from the case of an infinite crystal. Namely AHE on a finite lattice behaves in a similar way to a local Chern marker: its sum over all lattice sites $i,j$ is equal to zero, its bulk value is constant for a crystalline lattice and equal to the value it would have in a corresponding infinite lattice, and its value at the edges has opposite sign and large magnitude so that its sum over all sites near the eges exactly cancels out the sum over all sites in the bulk. At first sight, this may seem paradoxal, but there is a simple argument for it. Indeed it is well-known that on an infinite lattice, AHE is proportional to an integer topological invariant, and phases with different topological invariants cannot be smoothly transformed into each other, but are separated by phase transitions. However, it is also well-known that phase transitions do not exist on a finite lattice, but appear only in thermodynamic limit. This means that topologically distinct phases cannot exist on a finite lattice! In other words, any phase on a finite lattice is topologically equivalent to trivial. For this reason, if one sums the local Chern marker (3) over all lattice sites $r_{i}$, or one sums AHE (II.2.3) over all sites $i,j$, one obtains exactly zero.
@@
Figure 10: 10(a) Hall conductivity $\sigma_{i}$ (in units of $e^{2}/\hbar^{2}$) at each site $i$ plotted for lattices with $-L<x,y<L$ at various values of $m$, $t_{inter}$. One can see that in the bulk of a topological phase, it is approximately equal to $0.5$ Chern invariant divided by the number of sites in a unit cell in each layer, and in the non-topological phase, it is equal to zero. In the multifractal phase, it fluctuates within the lattice. 10(b) Averaged $\sigma_{i}$ summed over the four sublattice degrees of freedom ($A/B$ and top/bottom) and its standard deviation taken over the inner part of the lattice with $-L/2<x,y<L/2$. It remains constant and approximately equal to the bulk Chern invariant in the weakly coupled phase, but starts fluctuating at large $t_{inter}$.
@@ III Discussion
In this work, we have studied topological properties of quasicrystalline 30 degrees twisted bilayer honeycomb lattice. We started from the case of two uncoupled layers each of which forms Haldane model, and tracked its evolution once the interlayer coupling is increased. We found that at small interlayer coupling, two layers of crystalline topological insulators become a quasicrystalline topological insulator. The latter fact is not obvious because topological properties of Haldane model are characterized by Chern invariant defined in momentum space, which is in turn well-defined only in the presence of crystalline translational symmetry. In this regard, one could be concerned whether translational symmetry breaking might break topological protection, but we established that at small, but non-zero interlayer coupling, this is not the case: our bilayer has exactly the same topological properties as a superposition of two monolayers. Specifically, it still has edge states, has exactly the same value of topological entanglement entropy and the same behavior of local Chern marker and anomalous Hall conductivity.
As the interlayer coupling reaches a critical value, the system undergoes a ’hidden’ phase transition which manifests itself only through the behavior of local Chern markers and anomalous Hall conductivity. As the interlayer coupling increases further, our system undergoes a phase transition into a gapless phase, whose eigenstates are multifractal. Rigorous characterization of such a gapless phase’s topological properties is challenging, but one can just qualitatively say that its topology starts ’deteriorating’. Specifically, in the weakly coupled phases, the local Chern marker and anomalous Hall conductivity is constant in the bulk, but have a value of opposite sign in the boundary, whereas in the strongly coupled phase, these quantities significantly fluctuate within the bulk, but do not have any special features at the boundary. In fact, one may consider an analogy between our model and a disordered topological insulator. At weak disorder, the system has exactly the same topological properties as the topological insulator without disorder. However, once the disorder becomes strong, fluctuations of the disorder potential effectively behave like edges themselves, and this fact qualitatively explains that the bulk becomes indistinguishable from the physical edges of the system.
@@
It would be of interest to realize our model in experiments, but it is challenging because typically (e.g. in twisted bilayer graphene), at large twist angles, interlayer coupling is very weak. However, one may still try to make it larger by applying high pressure between the layer. More plausibly, one might try to realize our model using ultracold atoms (see e.g. [26]), which have been proved to be a powerful tool to realize various physical models.
@@ IV Acknowledgements
The author would like to thank Institute of Basic Science (Daejeon, S. Korea), in which this project started and Profs. Moon Jip Park, Kyong Min Kim, Sergej Flach, Sergey Syzranov, Predrag Nikolic, Renat Sabirianov, Luis Santos, Ivan Khaymovich for helpful discussions. Financial support by the National Science Foundation through EPSCoR RII Track-1: Emergent Quantum Materials and Technologies (EQUATE), Award OIA-2044049 is acknowledged.
@@ Appendix A Calculation of the entanglement entropy
$C_{ij}=\sum\limits_{filled}\psi(\vec{r}_{i})\psi^{\dagger}(\vec{r}_{j}).$
After that we select the components whose coordinates $\vec{r}_{i,j}$ are inside the subsystem whose entanglement entropy we are interested in and thus obtain ’reduced correlation matrix’ $C_{A,ij}$. The entanglement entropy can be expressed in terms of its eigenvalues $\zeta$ as
@@ Appendix B Kubo formula for anomalous Hall effect
$\mathcal{Z}=\int D\psi D\bar{\psi}e^{\sum\limits_{w,i}iwc^{\dagger}_{i,w}c_{i,w}-\sum\limits_{w,i,j}c^{\dagger}_{i,w}H_{ij}c_{j,w}}.$
@@
$\vec{J}_{ij,\Omega}=ic^{\dagger}_{i,w+\Omega}H_{ij}(\vec{r}_{j}-\vec{r}_{i})c_{j,w}$
@@
$J_{x,ij,\Omega}=\langle J_{x,ij,\Omega}\sum\limits_{k.l}J_{y,kl,-\Omega}\rangle A_{y,kl,\Omega}.$
@@
$G_{ij,w}=\sum\limits_{n}\frac{\psi_{i}\psi^{\dagger}_{j}}{iw-E_{n}}.$
After we sum over Matsubara frequencies, assume zero temperature and take the limit of zero external frequency, we can obtain an answer for $\sigma^{(Hall)}_{ij}$, which is given by the Eq. (II.2.3).
Finally, let us look at the sum $\sum\limits_{ij}\sigma^{(Hall)}_{ij}$ over all lattice sites $i,j$. We can simplify the Eq. (II.2.3) by using explicitly the fact that $\psi$ are eigenvectors, namely by applying an identity
@@
$\sum\limits_{i,j}\sigma^{(Hall)}_{ij}=\frac{1}{2}\sum\limits_{f,e}\sum\limits_{i,k}$
×$\times\left\{\psi^{\dagger}_{e,i}x_{i}\psi_{f,i}\cdot\psi^{\dagger}_{f,k}y_{k}\psi_{e,k}-\psi^{\dagger}_{f,i}H_{ij}x_{i}\psi_{e,i}\cdot\psi^{\dagger}_{e,k}y_{k}\psi_{f,k}\right.$
@@ Appendix C Anomalous Hall conductivity on an infinite crystalline lattice
The goal of this section is to show that the expression for Hall conductivity (II.2.3) from the main text in the case of an infinite crystalline lattice is indeed equivalent to a well-known expression in terms of Berry curvature. Here, for clarity we separate lattice indices $i,j....$ into pairs $(i,\alpha),(j,\beta),...$, which refer to unit cells and sublattice degrees of freedom respectively.
@@
$E\phi_{\alpha,k}=\sum\limits_{\beta}H_{\alpha\beta}(k)\phi_{\beta,k}$
@@
$H_{\alpha\beta}(k)=\sum\limits_{j,\beta}e^{i\vec{k}(\vec{x}_{j}-\vec{x}_{i})}H_{i,j\atop\alpha,\beta}.$
@@
$H_{i,j\atop\alpha,\beta}=\frac{1}{N}\sum\limits_{p}H_{\alpha\beta}(p)e^{-i\vec{p}(\vec{x}_{j}-\vec{x}_{i})}.$
After substituting the above expression, as well as expressions for the eigenstates in the form of plane waves (8) into the main expression for anomalous Hall conductivity (II.2.3), one can obtain the lattice version of TKNN formula
@@
×$\times\left\{\phi^{\dagger}_{e,\alpha}\frac{\partial H_{\alpha\beta}}{\partial k_{x}}\phi_{f,\beta}\cdot\phi^{\dagger}_{f,\gamma}\frac{\partial H_{\gamma\delta}(k)}{\partial k_{y}}\phi_{e,\delta}\right.$
@@
Notice that in the left side of this equation we have sum $\sum\limits_{j}\sigma_{ij}$ of conductivities through all bonds adjacent to a site $i$ the quantity we discuss in the main text. In addition, to obtain the physical answer we have to sum over all sublattice degrees of freedom, i.e. over the indices $\alpha,\beta$. We also note that in the Eq. (C) we still assume that the lattice is discreet, but the momentum is continuum because the lattice is infinite. Hence it is straightforward to replace the momentum summation with an integration
$\frac{1}{N}\sum\limits_{k}\to S\int\frac{d^{2}k}{(2\pi)^{2}},$

### v3

@@ II.3.1 Topological entanglement entropy
Figure 12: A scheme of partitioning the system, which we use to compute the topological contribution $\gamma$ to the entanglement entropy. Each of the circular curves has a radius of $L_{x}/2$. We compute entanglement entropy for each of the figures and then use this substraction scheme to isolate the contribution from the topological term $\gamma$.
We plot the results for $\gamma$ as functions of $m$, $t_{inter}$ on the Figs. 13(a) (we assume $E_{F}=0$). We can see that in the ’weakly coupled topological’ phase, $\gamma$ is a non-zero constant, which in turn confirms that it is a topological quasicrystal. In a similar way, in the ’weakly coupled non-topological’ phase, $\gamma$ is equal to zero, which means that the system is non-topological. Once the gap closes, the topological entanglement entropy becomes ill-defined.
We also compute $\gamma$ for large $t_{inter}$ at $m=0$, assuming that the Fermi level $E_{F}$ lies within the energy gap, which we previously found on the Figs 4(a), 4(b), 4(c). We obtain that at large system size, $\gamma$ becomes close to zero (see Fig. 14). This suggests that our ‘strongly coupled’ gapped system is not topological.
@@
Figure 13: Topological entanglement entropy $\gamma$ plotted against $t_{inter}$ (13(a)) for $m=0,2$ and against $m$ (13(b)) for $t_{inter}=0.0,1.0,10.0$. Here we assume $E_{F}=0$. It is equal to a non-zero constant in the topological gapped phase and zero in the non-topological phase
Figure 14: Topological entanglement entropy $\gamma$ plotted against $t_{inter}$ for $m=0$, in the case when the Fermi level lies within the bulk energy gap of the ‘strongly coupled’ system. We can see that with increasing system size $L_{x}$, it approaches zero.
II.3.2 Local Chern marker
@@
Figure 15: 15(a) Local Chern marker plotted for lattices with L<x,y<L-L<x,y<L at various values of $m$, $t_{inter}$. One can see that in the bulk of a topological phase, it is approximately equal to Chern invariant, and in the non-topological phase, it is equal to zero. In the multifractal phase, it fluctuates within the lattice. 15(b) Averaged local Chern marker $C_{ave}$ and its standard deviation $\Delta C$ taken over the inner part of the lattice with L/2<x,y<L/2-L/2<x,y<L/2. It remains constant and approximately equal to the bulk Chern invariant in the weakly coupled phase, but starts fluctuating at large $t_{inter}$.
II.3.3 Anomalous Hall conductivity
@@
$\times\left\{\psi^{\dagger}_{e,i}H_{ij}(x_{j}-x_{i})\psi_{f,j}\cdot\psi^{\dagger}_{f,k}H_{kl}(y_{l}-y_{k})\psi_{e,l}\right$.
@@
Let us emphasize once again, that in the Eq. (II.3.3), $\sigma_{ij}$ describes electric current through a bond between the lattice sites $i,j$ (we assumed that the electric field is uniform). However, one can sum electric currents through all bonds next to a given lattice site $i$ and consider $\sigma_{i}=\sum\limits_{j}\sigma_{ij}$. For a crystalline system, one can take a step further and split the lattice site index $i$ into two indices $i^{\prime},\alpha$ numerating unit cells and sublattice degrees of freedom respectively. The well-known integer value of the anomalous Hall conductivity is obtained after an additional summation over $\alpha$. A simple explanation of the latter is as follows: for an infinite crystalline system one may transform any dependency over unit cell index $i^{\prime}$ into momentum representation, and the well-known integer AHE occurs at zero momentum. However, its expression in terms of Berry curvature still contains summation over sublattice degrees of freedom. Nevertheless, on a honeycomb lattice, the two sublattice degrees of freedom are equivalent to each other, and therefore the AHE for each of them would be the same, i.e. half-integer.
More interestingly, properties of AHE obtained from the Eq. (II.3.3) on a finite lattice are fundamentally different from the case of an infinite crystal. Namely AHE on a finite lattice behaves in a similar way to a local Chern marker: its sum over all lattice sites $i,j$ is equal to zero, its bulk value is constant for a crystalline lattice and equal to the value it would have in a corresponding infinite lattice, and its value at the edges has opposite sign and large magnitude so that its sum over all sites near the eges exactly cancels out the sum over all sites in the bulk. At first sight, this may seem paradoxal, but there is a simple argument for it. Indeed it is well-known that on an infinite lattice, AHE is proportional to an integer topological invariant, and phases with different topological invariants cannot be smoothly transformed into each other, but are separated by phase transitions. However, it is also well-known that phase transitions do not exist on a finite lattice, but appear only in thermodynamic limit. This means that topologically distinct phases cannot exist on a finite lattice! In other words, any phase on a finite lattice is topologically equivalent to trivial. For this reason, if one sums the local Chern marker (7) over all lattice sites $r_{i}$, or one sums AHE (II.3.3) over all sites $i,j$, one obtains exactly zero.
@@
Figure 16: 16(a) Hall conductivity $\sigma_{i}$ (in units of $e^{2}/\hbar^{2}$) at each site $i$ plotted for lattices with L<x,y<L-L<x,y<L at various values of $m$, $t_{inter}$. One can see that in the bulk of a topological phase, it is approximately equal to $0.5$ - Chern invariant divided by the number of sites in a unit cell in each layer, and in the non-topological phase, it is equal to zero. In the multifractal phase, it fluctuates within the lattice. 16(b) Averaged $\sigma_{i}$ summed over the four sublattice degrees of freedom ($A/B$ and top/bottom) and its standard deviation taken over the inner part of the lattice with L/2<x,y<L/2-L/2<x,y<L/2. It remains constant and approximately equal to the bulk Chern invariant in the weakly coupled phase, but starts fluctuating at large $t_{inter}$.
@@ III Discussion
In this work, we have studied topological properties of quasicrystalline 30 degrees twisted bilayer honeycomb lattice. We started from the case of two uncoupled layers each of which forms Haldane model, and tracked its evolution with the increasing interlayer coupling. We found that at small interlayer coupling, our system exactly retains topological properties of the two uncouped layers: bulk gap and topological edge states persist, topological entanglement entropy remains constant, local Chern marker and anomalous Hall conductivity behave in exactly the same way as in the case of two uncoupled layers. This fact is not obvious because topological properties of Haldane model are characterized by Chern invariant defined in momentum space, which is in turn well-defined only in the presence of crystalline translational symmetry. In this regard, one could be concerned whether translational symmetry breaking might break topological protection, but we established that at small, but non-zero interlayer coupling, this is not the case: our bilayer has exactly the same topological properties as a superposition of two monolayers. In this sense, we demonstrated that two weakly coupled copies Chern insulators can be viewed as a quasicystalline topological insulator. In other words, we proposed to create a quasicystalline topological insulator by smoothly evolving crystalline topological insulator and breaking translational symmetry.
We found that as the interlayer coupling increases, our system undergoes a phase transition into a gapless phase, and its topological edge states disappear. As the interlayer coupling increases even further, another bulk gap may open. We found that such a stronly coupled phase is overall multifractal, but it possesses multiple states localized at various locations of the lattice. For example, at $m=0$, our system possesses states localized at the center of the lattice, as well as its corners. However, we believe that these corner modes are not topological in nature because their energies are not tied to the bulk gap, and also because topological entanglement entropy of the strongly coupled gapped system is almost zero. More rigorous topological classification of quasicrystals is a challenging and unresolved task, which we leave for future work.
We believe that closing of the topological band gap and the disappearance of edge modes at large interlayer coupling is natural to expect. In fact, one may consider an analogy between our model and a disordered topological insulator. At weak disorder, the system has exactly the same topological properties as the topological insulator without disorder. However, once the disorder becomes strong, fluctuations of the disorder potential effectively behave like edges themselves, and this fact qualitatively explains the gap closing and implies that the bulk becomes indistinguishable from the physical edges of the system.
@@
It would be of interest to realize our model in experiments, but it is challenging because typically (e.g. in twisted bilayer graphene), at large twist angles, interlayer coupling is very weak. However, we expect that our system should host qualitatively similar properties not only at $30^{\circ}$ twist, but at any incommensurate twist angle. Moreover, one could create quasiperiodicity not only by an incommensurate twist angle, but also by stacking together two monolayers with different lattice constants. The latter is plausible to achieve considering the fact that the family of 2D materials by now is very rich (e.g. graphene, transition metal dichalcocenides etc). One may also try to change interlayer spacing by applying high pressure between the layers. Finally, we would like to point out the possibility to realize our model using ultracold atoms (see e.g. [28]), which have been proved to be a powerful tool to realize various physical models.
@@ IV Acknowledgements
The author would like to thank Institute of Basic Science (Daejeon, S. Korea), in which this project started and Profs. Moon Jip Park, Kyong Min Kim, Sergej Flach, Sergey Syzranov, Predrag Nikolic, Renat Sabirianov, Luis Santos, Ivan Khaymovich, Niccolo Traverso Ziani, Matteo Carrega for helpful discussions. Financial support by the National Science Foundation through EPSCoR RII Track-1: Emergent Quantum Materials and Technologies (EQUATE), Award OIA-2044049 is acknowledged.
@@ Appendix A Calculation of the entanglement entropy
$C_{ij}=\sum\limits_{filled}\psi(\vec{r}_{i})\psi^{\dagger}(\vec{r}_{j}).$ (10)
After that we select the components whose coordinates $\vec{r}_{i,j}$ are inside the subsystem whose entanglement entropy we are interested in and thus obtain ’reduced correlation matrix’ $C_{A,ij}$. The entanglement entropy can be expressed in terms of its eigenvalues $\zeta_{m}$ as
@@ Appendix B Kubo formula for anomalous Hall effect
$\mathcal{Z}=\int D\psi D\bar{\psi}e^{\sum\limits_{w,i}iwc^{\dagger}_{i,w}c_{i,w}-\sum\limits_{w,i,j}c^{\dagger}_{i,w}H_{ij}c_{j,w}}.$ (12)
@@
$\vec{J}_{ij,\Omega}=ic^{\dagger}_{i,w+\Omega}H_{ij}(\vec{r}_{j}-\vec{r}_{i})c_{j,w}$ (14)
@@
$J_{x,ij,\Omega}=\langle J_{x,ij,\Omega}\sum\limits_{k.l}J_{y,kl,-\Omega}\rangle A_{y,kl,\Omega}.$ (15)
@@
$G_{ij,w}=\sum\limits_{n}\frac{\psi_{i}\psi^{\dagger}_{j}}{iw-E_{n}}.$ (17)
After we sum over Matsubara frequencies, assume zero temperature and take the limit of zero external frequency, we can obtain an answer for $\sigma^{(Hall)}_{ij}$, which is given by the Eq. (II.3.3).
Finally, let us look at the sum $\sum\limits_{ij}\sigma^{(Hall)}_{ij}$ over all lattice sites $i,j$. We can simplify the Eq. (II.3.3) by using explicitly the fact that $\psi$ are eigenvectors, namely by applying an identity
@@
$\sum\limits_{i,j}\sigma^{(Hall)}_{ij}=\frac{1}{2}\sum\limits_{f,e}\sum\limits_{i,k}$ (19)
$\times\left\{\psi^{\dagger}_{e,i}x_{i}\psi_{f,i}\cdot\psi^{\dagger}_{f,k}y_{k}\psi_{e,k}-\psi^{\dagger}_{f,i}H_{ij}x_{i}\psi_{e,i}\cdot\psi^{\dagger}_{e,k}y_{k}\psi_{f,k}\right$.
@@ Appendix C Anomalous Hall conductivity on an infinite crystalline lattice
The goal of this section is to show that the expression for Hall conductivity (II.3.3) from the main text in the case of an infinite crystalline lattice is indeed equivalent to a well-known expression in terms of Berry curvature. Here, for clarity we separate lattice indices $i,j.$... into pairs $(i,\alpha),(j,\beta),$..., which refer to unit cells and sublattice degrees of freedom respectively.
@@
$E\phi_{\alpha,k}=\sum\limits_{\beta}H_{\alpha\beta}(k)\phi_{\beta,k}$ (23)
@@
$H_{\alpha\beta}(k)=\sum\limits_{j,\beta}e^{i\vec{k}(\vec{x}_{j}-\vec{x}_{i})}H_{i,j\atop\alpha,\beta}.$ (24)
@@
$H_{i,j\atop\alpha,\beta}=\frac{1}{N}\sum\limits_{p}H_{\alpha\beta}(p)e^{-i\vec{p}(\vec{x}_{j}-\vec{x}_{i})}.$ (25)
After substituting the above expression, as well as expressions for the eigenstates in the form of plane waves (22) into the main expression for anomalous Hall conductivity (II.3.3), one can obtain the lattice version of TKNN formula
@@
$\times\left\{\phi^{\dagger}_{e,\alpha}\frac{\partial H_{\alpha\beta}}{\partial k_{x}}\phi_{f,\beta}\cdot\phi^{\dagger}_{f,\gamma}\frac{\partial H_{\gamma\delta}(k)}{\partial k_{y}}\phi_{e,\delta}\right$.
@@
Notice that in the left side of this equation we have sum $\sum\limits_{j}\sigma_{ij}$ of conductivities through all bonds adjacent to a site $i$ - the quantity we discuss in the main text. In addition, to obtain the physical answer we have to sum over all sublattice degrees of freedom, i.e. over the indices $\alpha,\beta$. We also note that in the Eq. (C) we still assume that the lattice is discreet, but the momentum is continuum because the lattice is infinite. Hence it is straightforward to replace the momentum summation with an integration
$\frac{1}{N}\sum\limits_{k}\to S\int\frac{d^{2}k}{(2\pi)^{2}},$ (27)